OpenStudy (anonymous):

for the joint p.d.f: f(x,y)=c(x^2+y) ;0<=y<=1-x^2 0 ; elsewhere determine the value of constant c

OpenStudy (anonymous):

\[f(x,y)= c(x^{2}+y); 0\le y \le1-x^{2}\]

OpenStudy (zarkon):

what are the bounds on x?

OpenStudy (anonymous):

not given

OpenStudy (anonymous):

just the above condition

OpenStudy (zarkon):

I assume then at \(-1\le x\le 1\)

OpenStudy (anonymous):